OpenStudy (anonymous):

What is the product in scientific notation? (3.5x10^-5)(3x10-10)

OpenStudy (welshfella):

3.5 * 3 = ? 10^-5 * 10^-10 = 10^(-5-10) = ?

OpenStudy (anonymous):

3.5*3=10.5

OpenStudy (anonymous):

I Just got confused

OpenStudy (welshfella):

for the second part work out -5-10 and put it as an exponent over 10

OpenStudy (welshfella):

I've assumed that 10-10 is 10^-10

OpenStudy (welshfella):

is it that or is it 3x10 minus 10?